Use the elimination method to solve the given system of equations. Please show all work
2a - 4b = 9;
3a + 4b = 1

Answer by jim_thompson5910(35256)

Hint: add the two equations to get


2a - 4b = 9
3a + 4b = 1
------------
5a+0b = 10


This means 5a = 10. Solve this equation for 'a'. Then use this solution for 'a' to plug it back into an equation with a,b in it solve for b.